How many Mg ions are present in 3.00 moles of MgCl2?
1
Expert's answer
2020-02-14T05:47:38-0500

Task:

How many Mg ions are present in 3.00 moles of MgCl2?


Solution:

In water MgCl2 dissociates to ions: a cation of magnesium (Mg2+) and two anions of cloride (Cl-).

MgCl2 = Mg2+ + 2Cl-

Hence, n(MgCl2) = n(Mg2+) = 3.00 moles (according to the reaction equation).

n = N / Na, where Na - Avogadro's number = 6.022*1023 mol-1.

Then,

N(Mg2+) = n(Mg2+) * Na = 3.00 * 6.022*1023 = 18.066*1023 = 1.81*1024


Answer: 1.81*1024 Mg2+ ions are present in 3.00 moles of MgCl2.